Hal purchases a car for $65,000, makes a down payment of 30%, and finances the rest with a 3-year car loan at an annual interest rate of 6.2% compounded monthly. What is the amount of his monthly loan payment?
Answer by addingup(3677)

65000-(65000*0.3) = 45500
He finances 45500 over 36 months at 6.2% compounded monthly and 0.062/12 = 0.00517
---------------------------------------
Monthly Pmt = (0.00517*45500)/(1-(1+0.00517)^-36)
= 235.235/(1-0.8306) = 1388.64